40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1329     Entry time: Fri Feb 20 03:52:23 2009
Author: Yoichi 
Type: Update 
Category: Locking 
Subject: Locking Tonight 
Yoichi, Peter

Tonight, we had a problem with the DD hand off.
It failed when the RG filters of MICH for the bounce-roll modes are engaged.
The reason for the failure was that the MICH UGF was too low (~10Hz).
As in the Peter's elog entry, we found that the AS PDs are mis-centered.
Even after we fixed the centering, the MICH UGF was still too low. So we increased the MICH feedback gain by a factor of 10.
The reason for the gain decrease is unknown. It seems almost like the BS coils get weaker.
I checked the UGF of the BS OL loops. These are around 4Hz, so fine. We should check the HWP on the AP table tomorrow.

After the DD hand-off goes ok, the switching of DARM signal from DC to RF failed.
I found that the gain and the polarity of the RF signal were wrong.
AS166 is one of the PDs we found mis-centered (and re-centered). But how can you flip the sign of the signal ?

After this, the cm_step script goes until the activation of the moving zero, but fails when the arm power is increased to 0.7.
Also the ontoMCL script succeeds only 50% of the time.
ELOG V3.1.3-